OpenStudy (momo):

Find the slope of the line passing through the points (1, 4) and (7, 2).

OpenStudy (anonymous):

m = (2 - 4) / (7 - 1) = -2 / 6 = -1 / 3

OpenStudy (momo):

thanks ^-^